A total of
596
tickets were sold for the school play. They were either adult tickets or student tickets. There were
54
fewer student tickets sold than adult tickets. How many adult tickets were sold?

Found 2 solutions by math_helper, stanbon:
Answer by math_helper(2461)

S = number of student tickets
A = number of adult tickets
S+A = 596 (1)
S = A-54 (2)

Subs A-54 for S from (2) , into (1):
(A-54)+A = 596
2A - 54 = 596

Solve for A then plug in that value into either eqn (1) or (2) to get S.

Answer by stanbon(75887)

A total of
596
tickets were sold for the school play. They were either adult tickets or student tickets. There were
54
fewer student tickets sold than adult tickets. How many adult tickets were sold?
--------------
Quantity Eq:: a + s = 596 tickets
Quantity Eq:: s = a - 54
------------------------------
Substitute for "s" and solve for "a"::
a + a-54 = 596
------------------
2a = 650
adult tickets sold = 325
student tickets sold = 325-54 = 271
------------------
Cheers,
Stan H.
-------------